A bowl of water has a temperature of 50◦C. It is put into a refrigerator where the temperature is 5◦C. After 0.5 hours, the water is stirred and its temperature is measured to be 20◦C. It is then left to cool further, take t in Newton’s law of cooling to be measured in minutes. Use Newton’s law of cooling to predict when the temperature will be 10◦C.
1
Expert's answer
2020-06-30T15:29:16-0400

"T(t)=T_s+(T_o-T_s)e^{-kt}\\
T(t) \text{ --- temperature of the object at time t}\\
T_s \text{ --- temperature of surrounding}\\
T_o \text{ --- temperature of the object at time } t=0\\
k=\text{const}\\
\text{A bowl of water has a temperature of 50◦C. It is put}\\
\text{into a refrigerator where the temperature is 5◦C}.\\
\text{After 0.5 hours, the water is stirred and its temperature}\\
\text{is measured to be 20◦C. So we have:}\\
20=5+(50-5)e^{-0.5k}\\
e^{-0.5k}=\frac{1}{3}\\
-(0.5)k=\ln\frac{1}{3}\\
k=\ln 9\\
\text{It is then left to cool further. We should predict when}\\
\text{the temperature will be 10◦C:}\\
10=5+(20-5)e^{-(\ln 9) t}\\
e^{-(\ln 9) t}=\frac{1}{3}\\
t=0.5\\
\text{The temperature will be 10◦C in 30 minutes}\\
\text{(after we put a bowl of water into a refrigerator the second time)}."